Scores on a test have a mean of 78.4 and 7 percent of the scores are above 90. The scores have a distribution that is approximately normal. Find the standard deviation. Round your answer to the nearest tenth, if necessary.
Answer #1

Normal distribution: P(X < A) = P(Z < (A - )/)

Mean, = 78.4

P(X > 90) = 0.07

P(X < 90) = 1 - 0.07 = 0.93

P(Z < (90 - 78.4)/) = 0.93

(90 - 78.4)/ = 1.48

= 7.8

Standard deviation of the distribution is 7.8